OpenStudy (anonymous):

radius of convergence: 2x^(2n)/n+1

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} 2x ^{2n}/(n+1)\]

OpenStudy (anonymous):

using the ratio test

OpenStudy (anonymous):

\[2x ^{2(n+1)}/((n+1)+1) \times (n+1)/(2x ^{2n})\]

OpenStudy (anonymous):

i am going to guess (-1,1)

OpenStudy (anonymous):

hmm... but i have a problem with the next step

OpenStudy (anonymous):

wait guys..

OpenStudy (anonymous):

That's the interval of convergence. The radius of convergence would be 1 in that case.

OpenStudy (anonymous):

yes true

OpenStudy (anonymous):

this one the first fraction \[2x ^{2n+2}/n+2\]

OpenStudy (anonymous):

yes so compute \[\frac{x^{2n+2}}{n+2}\times \frac{n+1}{x^{2n}}\]

OpenStudy (anonymous):

how can i rewrite this: like? \[2x ^{2n} * 4x ^{2} / (n+2)\] or like \[2x ^{2n} * 2x ^{2}\]

OpenStudy (anonymous):

how do you insert the fraction with the equation?

OpenStudy (anonymous):

the ratio is the one i wrote. it is literally \[\frac{a_{n+1}}{an}\] but for practical purposes you almost always have a ratio to begin with so you multiply by the reciprocal as you would normally

OpenStudy (anonymous):

\frac{a}{b}

OpenStudy (anonymous):

\[\frac{2x ^{2n}4x ^{2}}{n+2} \times \frac{n+1}{2x ^{2n}}\] or \[\frac{2x ^{2n}2x ^{2}}{n+2} \times \frac{n+1}{2x ^{2n}}\]

OpenStudy (anonymous):

im confussed with the \[4x^{2}\] and \[2x^{2}\]

OpenStudy (anonymous):

ignore the 2 take it right out front of the sum. it is a red herring

OpenStudy (anonymous):

or just cancel at that step

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

if i cancel that step i dont know how to go further

OpenStudy (anonymous):

ii think that 4 is a mistake

OpenStudy (anonymous):

replace n by n + 1 in your expression. it does not effect the "2"

OpenStudy (anonymous):

could you write out what you are saying?

OpenStudy (anonymous):

i dont understand it anymore..

OpenStudy (anonymous):

1 thing i know is that the 2x^2n term dissapears

OpenStudy (anonymous):

yes ok \[a_n=\frac{2x^{2n}}{n+1}\] \[a_{n+1}=\frac{2x^{2(n+1)}}{n+1+1}=\frac{2x^{n+2}}{n+2}\]

OpenStudy (anonymous):

so there is no 4 in it. increase n by 1 is all your ratio is therefore \[\frac{2x^{2n+2}}{n+2}\times \frac{n+1}{2x^{2n}}\] \[=\frac{x^2(n+1)}{n+2}\]

OpenStudy (anonymous):

but isnt \[2x ^{2n+2} = 2x ^{2n}*2n ^{2}?\]

OpenStudy (anonymous):

oops forgot the x

OpenStudy (anonymous):

the n is a x

OpenStudy (anonymous):

owh wait i got it... the 2 is just a scalar :) nvm ^^ thank youu <3

OpenStudy (anonymous):

yw, and yes, the 2 you could safely ignore